So there are two similar triangles, The larger one has (A)side that measures 34 ft, and (B) side that measures 9 ft. The smaller triangle has an (A) side that measures 20 ft and a (B) side that measures Y.
Find the unknown length y in the pair of similar triangles.
(Simplify your answer. Type an integer, fraction, or mixed number.)
Can anyone help. I'm stuck on this problem.
(Sorry, but I dont know how to draw triangles on this)

Answer by edjones(7569)

y/9 = 20/34
y = 180/34 = 90/17 ft